OpenStudy (anonymous):

Evaluate 10 + 4(3 + 2) + 5 + (12 ÷ 6).

OpenStudy (anonymous):

10 + 4(3 + 2) + 5 + (12 ÷ 6) Using BODMAS, 10 + 4(3 + 2) + 5 + (12 ÷ 6) = 10 + 4(5) + 5 + 2 = 10 + 20 + 5 + 2 = 37

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You are welcome